# AlphaNet: Scaling Up Local Frame-based Neural Interatomic Potential

### v2

@@ Abstract
We present AlphaNet, a local frame-based equivariant model designed to achieve both accurate and efficient simulations for atomistic systems. Recently, machine learning force fields (MLFFs) have gained prominence in molecular dynamics simulations due to their advantageous efficiency-accuracy balance compared to classical force fields and quantum mechanical calculations, alongside their transferability across various systems. Despite the advancements in improving model accuracy, the efficiency and scalability of MLFFs remain significant obstacles in practical applications. AlphaNet enhances computational efficiency and accuracy by leveraging the local geometric structures of atomic environments through the construction of equivariant local frames and learnable frame transitions. We substantiate the efficacy of AlphaNet across diverse datasets, including defected graphene, formate decomposition, zeolites, and surface reactions. AlphaNet consistently surpasses well-established models, such as NequIP and DeepPot, in terms of both energy and force prediction accuracy. Notably, AlphaNet offers one of the best trade-offs between computational efficiency and accuracy among existing models. Moreover, AlphaNet exhibits scalability across a broad spectrum of system and dataset sizes, affirming its versatility. Our code and data is available at https://github.com/zmyybc/AlphaNet.
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com and haixiao@tsinghua.edu.cn.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [30, 4, 24]. While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps). As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [8, 25, 19]. However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
@@
The advent of machine learning techniques has introduced a promising solution to this dilemma. By training models on data derived from first-principles calculations, machine learning force fields (MLFFs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
@@
As interest in applying MLFFs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 [43], is a large-scale pre-trained model with improved attention architecture over DeepPot [40]. Similarly, JMP [35] leverages diverse molecular systems of different types as a joint pre-training model. MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of MLFF-driven atomistic simulations [14, 2, 33, 36, 9].
@@
Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important. Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection. Thus, equivariant models are developed to respect these symmetries  [32, 38], and most of these expressive models by date [2, 1, 23] are based on the spherical harmonics [17], but calculating the tensor product of irreducible representations imposes expensive computational overheads. While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [20, 10, 41, 13], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
@@
In this paper, we propose a local-frame-based atomistic equivariant model, named AlphaNet, for accurate yet highly efficient atomistic simulations. Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art MLFF models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, to surface reactions.
2 Experiment
2.1 Dataset
The Defected Bilayer Graphene Dataset consists of reference structures specifically designed to train and validate machine learning potentials (MLPs). It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers). The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures. Additionally, snapshot configurations from MD simulations at different temperatures were included. The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively. Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.  [42]
The Formate Decomposition on Cu Dataset consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond. The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase. Nudged elastic band (NEB) method [18] was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations. These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory. AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.  [2]
The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation. Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability. The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) [5] simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) [21, 22]. For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration. The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
The OC2M Dataset is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials [6]. This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both). Each of them contains around 1 million structures. It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments. These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures. Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration. This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
The MPTraj Dataset and Matbench Discovery Test. MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [9]. Matbench Discovery [31] is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [39]. In this case we train our model on the MPTraj dataset and test it on the WBM test set.
Figure 1: Illustration of Experimental Systems: a) The Formate Decomposition on Cu Dataset. Cu atoms are depicted in brown, C atoms in grey, O atoms in red, and H atoms in white. b) The Zeolite Dataset. Li atoms are shown in purple, Si atoms in grey, Al atoms in orange, and O atoms in red. c) The Defected Bilayer Graphene Dataset.
2.2 Results
2.2.1 Accuracy
Our model demonstrates excellent precision across various experimental datasets. On datasets such as formate decomposition and defected graphene, as detailed in Table 1. We selected NequIP [2] as our baseline model due to its demonstrated superior performance in the research associated with the dataset. AlphaNet consistently outperforms the baseline model in terms of both energy and force accuracy.
@@ 2.2.1 Accuracy
The formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface. For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Å for force and 0.23 meV/atom for energy, compared to NequIP’s 47.3 meV/Å and 0.50 meV/atom, respectively. These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules. This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
@@
Another challenge for machine learning force fields (MLFFs) is accurately modeling the sliding layer effect in materials like defected graphene, where interlayer interactions are critical. On the defected graphene test set, AlphaNet achieves a force MAE of 32.0 meV/Å and an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Å for force MAE and 1.9 meV/atom for energy MAE. Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections. The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 3).
This ability to model both metallic and covalent bonding scenarios, as well as interlayer interactions, makes AlphaNet a versatile and powerful tool for materials design, particularly in fields requiring high accuracy for large-scale systems. These results suggest that AlphaNet can be effectively applied to a wide variety of systems, ranging from catalytic surfaces to layered materials, where precision in both energy and force calculations is critical for predicting material behavior.
2.2.2 Scaling
In the context of machine learning force field (MLFF) models, performance across large and diverse datasets is crucial for validating model robustness and scalability. The zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations, provides an excellent benchmark to evaluate the accuracy of these models. Here, we compare the performance of Deep Potential [40] and AlphaNet using two key metrics: Energy MAE (meV/atom) and Force MAE (meV/Å). A lower mean absolute error (MAE) indicates higher predictive accuracy.
Table 5 shows the results of Deep Pot and AlphaNet across different zeolite configurations. Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions.
For instance, in the ABWopt configuration, AlphaNet achieves an Energy MAE of 41 meV, substantially lower than Deep Pot’s 90 meV. Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å. This pattern of superiority is consistent across most configurations. Notable results include the CANopt structure, where AlphaNet shows an exceptional Force MAE of 10 meV/Å compared to Deep Pot’s 90 meV/Å, indicating a marked improvement in force prediction accuracy. Additionally, in the FERopt structure, AlphaNet achieves an Energy MAE of 100 meV, significantly lower than Deep Pot’s 290 meV, and a Force MAE of 18 meV/Å, much better than Deep Pot’s 130 meV/Å.
The performance of the model in terms of Mean Absolute Error (MAE) for both energy (in meV) and force (in meV/Å) improves as the training data size increases from 5k to 80k samples. Specifically, as shown in Figure 2, the energy MAE decreases from 220 meV at 5k samples to 41 meV at 80k samples. Similarly, the force MAE decreases from 241 meV/Å at 5k samples to 54 meV/Å at 80k samples. This trend highlights the scalable performance of increasing the training dataset size in enhancing the accuracy of both energy and force predictions on the zeolite dataset.
(a) Energy MAE (meV) vs Data Size
(b) Force MAE (meV/Å) vs Data Size
Figure 2: MAE of Energy and Force predictions as a function of training data size on the zeolite dataset.
Even in cases where Deep Pot performs competitively, such as in the LOSopt configuration, where Deep Pot achieves an Energy MAE of 110 meV/atom compared to AlphaNet’s 130 meV, AlphaNet still maintains an advantage in Force MAE with a value of 46 meV/Å compared to Deep Pot’s 70 meV/Å. This suggests that while Deep Pot may occasionally perform well in energy prediction, AlphaNet is more consistent in providing better overall results, particularly in force predictions.
In terms of standout performance, AlphaNet demonstrates its predictive power in configurations like LTJopt, where it achieves an Energy MAE of just 1.2 meV compared to Deep Pot’s 70 meV, and in the GISopt structure, where it delivers a Force MAE of only 3.1 meV/Å, far surpassing Deep Pot’s 50 meV/Å.
These results indicate that AlphaNet is better equipped to handle complex zeolite configurations, providing both lower energy and force prediction errors across most datasets. This demonstrates its robustness and makes it a promising model for large-scale applications in materials design, where accuracy and scalability are critical.
The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research. In our current work, we focus on scaling our model to the OC2M S2EF task, with plans to extend training to the full OC20 dataset in future research. We compare the performance of our model against established benchmarks, including EquiformerV2 [23], EScAIP [29], eSCN [28], GemNet-OC [16], SchNet [33], and DimeNet++ [15].
Our comparisons include SchNet and DimeNet++ models trained on the full OC20 dataset ($\sim$134M data points), as well as EquiformerV2, EScAIP, eSCN, and GemNet-OC trained on the smaller 2M subset. Specifically, we utilize pretrained models of SchNet and DimeNet++ available at https://fair-chem.github.io/core/model_checkpoints.html, while results for the other models are sourced from [29]. All reported results are based on the validation set, as it is publicly accessible. The corresponding performance metrics are summarized in Table 2.
To manage training costs and ensure exposure to all elements in the dataset, we first train our model on the full OC training set for 10,000 steps before switching to the 2M subset. We note that this initial step does not provide additional benefits, as testing confirms no significant difference in performance on the validation ID set, which contains no additional elements. Our model, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV. This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit higher MAE values above 0.35 eV. This trend remains consistent across various test sets, highlighting the robustness of our model in energy prediction.
In terms of force prediction, while our model trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP, it remains competitive relative to models of similar scale trained on the full dataset. This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$. Table 3 provides an efficiency evaluation of different models. Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than models like EScAIP. It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for autograd, resulting in higher time and memory costs.
Speak to WBM test, AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4. With an F1 score of 0.789 and DAF score of 4.315, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints. Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$). This combination of efficiency and physics-aware design positions AlphaNet as a robust foundation for scalable materials modeling.
Given these promising results, we are confident that our model can be successfully scaled to the full OC20 dataset in future work, further enhancing its predictive accuracy for both energy and force calculations. This would allow us to continue improving the state-of-the-art in machine learning for catalysis research.
2.2.3 Speed
The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed. While increasing the number of parameters in an ML model can enhance its precision, this often comes at the cost of slower inference times. To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10. Additionally, we tested systems of varying sizes by expanding the unit cells. Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
We compared several graph-based models, including NequIP, MACE, SchNet, and PaiNN, against our AlphaNet. As illustrated in Figure 4 (a), AlphaNet demonstrates a distinct speed advantage over other MLFFs, exhibiting a relatively slow increase in computational time as system size grows. This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
Furthermore, we evaluated the inference speed of pretrained models on the Matbench Discovery leaderboard. To ensure fairness, we selected models that are compliant, open-source, and accessible without additional requests. We measured the average speed of energy prediction using the same API. The models included AlphaNet, ORB-v2 [26], SevenNet [27], GRACE [3], MACE, and CHGNet. Figure 4 (b) presents the results. Among these models, GRACE, which is based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based. Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-0 and SevenNet-l3i5). This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
Table 1: Performance of Models on Different Datasets
@@ 2.2.3 Speed
Formate Decomposition Force MAE (meV/Å) 47.3 45.5
Energy MAE (meV/atom) 0.50 0.23
Defected Graphene Force MAE (meV/Å) 60.2 32.0
Energy MAE (meV/atom) 1.9 1.7
Table 2: Results on the OC20 validation set
Model Parameters ↓ Energy MAE (eV) ↓ Force MAE (eV/Å) ↓
SchNet-full 9.1M 0.54 0.55
DimeNet++-full 10.1M 0.53 0.048
GemNet-OC 38M 0.29 0.026

### v3

# AlphaNet: Scaling Up Local Frame-based Neural Network Interatomic Potentials

@@ Abstract
Molecular dynamics simulations demand unprecedented accuracy and scalability to tackle grand challenges in energy materials, catalytic processes, and biomolecular design. To bridge this gap, we present AlphaNet, a local frame-based equivariant model that simultaneously advances computational efficiency and predictive precision for atomistic systems. By constructing equivariant local frames with learnable geometric transitions, AlphaNet encodes atomic environments with enhanced representational capacity, achieving state-of-the-art accuracy in energy and force predictions. Extensive benchmarks—spanning defected graphene, formate decomposition, inorganic bulks, and large-scale datasets (OC2M and Matbench Discovery) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes. The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for simulating multiscale phenomena, from catalyst dynamics to energy storage interfaces, with direct implications for accelerating the discovery of functional materials and complex molecular systems. Our code and data is available at https://github.com/zmyybc/AlphaNet.
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@gmail.com and haixiao@tsinghua.edu.cn.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems (Richards et al., 2016; Boero et al., 1998; Lindorff-Larsen et al., 2011). While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps). As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories (Jones and Chapman, 1924; Cornell et al., 1995; MacKerell et al., 1998). However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics (Iftimie et al., 2005).
@@
The advent of machine learning techniques has introduced a promising solution to this dilemma. By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
@@
As interest in applying NNIPs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 (Zhang et al., 2024a), is a large-scale pre-trained model with improved attention architecture over DeepPot (Wang et al., 2018). Similarly, JMP (Shoghi et al., 2024) leverages diverse molecular systems of different types as a joint pre-training model. MACE (Batatia et al., 2022), on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations (Qu and Krishnapriyan, 2024; Liao et al., 2024; Gasteiger et al., 2021; Batzner et al., 2022; Schütt et al., 2018; Smith et al., 2017; Deng et al., 2023).
@@
Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important. Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection. Thus, equivariant models are developed to respect these symmetries  (Satorras et al., 2021; Thomas et al., 2018), and most of these expressive models by date (Batzner et al., 2022; Batatia et al., 2022; Liao et al., 2024) are based on the spherical harmonics (Geiger and Smidt, 2022), but calculating the tensor product of irreducible representations imposes expensive computational overheads. While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global (Jumper et al., 2021; Du et al., 2022, 2023; Duval et al., 2023), and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
@@
In this paper, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations. Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, surface reactions and inorganic bulks.
2 Results
We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset size, and computational efficiency in terms of inference speed and memory usage. Initially, we validate atomic-level prediction accuracy across five diverse chemical systems, simultaneously demonstrating AlphaNet’s scalability concerning dataset and system sizes. Subsequently, we quantify computational efficiency by benchmarking both the memory footprint and inference costs of the model. Comprehensive details regarding datasets, comparative methodologies, and hyperparameter configurations are provided in Appendix B.
2.1 AlphaNet achieving high accuracy across all datasets.
@@ 2 Results
We first consider the formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface. For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP (Batzner et al., 2022) 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1. These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules. This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
@@
Another challenge for NNIP models is accurately modeling the sliding layer effect in materials like defected graphene, where inter-layer interactions are critical. On the defected graphene dataset, AlphaNet achieves a force MAE of 32.0 meV/Åand an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Åfor force MAE and 1.9 meV/atom for energy MAE. Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections. The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 1).
Table 1: Results on the Formate Decomposition and Defected Graphene datasets. Boldface indicates the best performance.
@@
Formate Decomposition Force MAE (meV/Å) $\downarrow$ 47.3 45.5
Energy MAE (meV/atom) $\downarrow$ 0.50 0.23
Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 32.0
Energy MAE (meV/atom) $\downarrow$ 1.9 1.7
(a) Parity plot for total energies obtained for the training (blue), validation (orange), and test (green) data sets.
(b) Parity plot for atomic forces obtained for the training (blue), validation (orange), and test (green) data sets.
(c) AlphaNet (middle) and DFT (left) sliding PES for bilayer graphene, and their difference (right).
Figure 1: Comparisons between AlphaNet predictions and DFT calculations on the Zeolite dataset.
The third dataset we consider is the zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations. Table 2 shows the results of Deep Pot (Wang et al., 2018) and AlphaNet across different zeolite configurations. Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions. For instance, in the ABWopt configuration, AlphaNet achieves an Energy MAE of 41 meV, substantially lower than Deep Pot’s 90 meV. Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å.
Table 2: Results on the Zeolite dataset. Boldface indicates the best performance. (Unit for energy is meV/atom and unit for force is meV/Å).
ABWopt BCTopt BPHopt CANopt
Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
Energy MAE $\downarrow$ 90 71 110 73 210 41 150 62
Force MAE $\downarrow$ 90 54 50 55 60 90 90 10
EDIopt FERopt GISopt JBWopt
Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
Energy MAE $\downarrow$ 40 5.4 290 100 60 56 150 18
Force MAE $\downarrow$ 50 59 130 18 50 3.1 70 50
LOSopt LTAopt LTJopt NATopt
Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
Energy MAE $\downarrow$ 110 130 150 90 70 1.2 210 47
Force MAE $\downarrow$ 70 46 64 60 110 60 110 9.9
PARopt PHIopt SODopt THOopt
Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
Energy MAE $\downarrow$ 90 6.5 60 16 200 9.5 160 70
Force MAE $\downarrow$ 70 3.5 120 50 110 21 60 47
Table 3: Results on the OC20 validation set. “full” denotes models trained on the full OC20 training set, others are trained on the 2M subset. $\lambda_{E}$ refers to the weight of the energy loss, the weight of the force loss $\lambda_{F}$ is set to $100$. Boldface indicates the best performance.
Model Parameters $\downarrow$ Energy MAE (eV) $\downarrow$ Force MAE (eV/Å) $\downarrow$
SchNet (full) 9.1M 0.54 0.55
DimeNet++ (full) 10.1M 0.53 0.048
GemNet 38M 0.29 0.026